Operation Summary for March 21, 2011 - March 23, 2011 |
|
Monday March 21 The day shift began with Operations monitoring the Recycler stash (513.4E10), with the Antiproton Source (Pbar) stacking (23.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE and MTest experiments taking beam, and with the Tevatron off for D16-2 magnet replacement . During the day shift, a new record for the Recycler stash (608.6E10) was set. Operations put F-Sector into access mode at 1:03 PM, beam was disabled to Pbar, MiniBooNE and MTest. MiniBooNE beam returned at 2:17 PM, beam remained off to Pbar and MTest. During the evening shift, Operations sent beam to MiniBooNE. Tuesday March 22 The midnight shift began with Operations monitoring the stash (546.9E10), and with MiniBooNE take beam. During the midnight shift, Operations established beam to MTest at 4:45 AM. Problems with Linac RF Station #1 (LRF1) keep beam off for the last 90 minutes of the shift. During the day shift, Linac experts had LRF1 back on by 8:30 AM, beam returned to MiniBooNe and MTest at this time. Problems with Main Injector RF caused about one hour of downtime. During the evening shift, Operations turned off MTest at 10:00 PM. Wednesday March 23 The midnight shift began with Operations monitoring the stash (522.8E10), and with MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4:47 AM. Operations disabled beam to MiniBooNE and MTest at 6:29 AM for Transfer Hall and F-Sector accesses (related to D16-2 magnet replacement). The Plans for Wednesday The plans for today are to finish Transfer Hall and F-Sector accesses and return beam to MiniBooNE and MTest. |
|
More Information Comments and Suggestions |
| last modified 03/23/2011 email Fermilab |
| Security, Privacy, Legal | |